You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121207 (28 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 6-7

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: M. Edwars
Instrument: LUCI1

Summary

Clear sky at the beginning of the observations.
At the beginning of the night the average seeing is 1.0".
Strong wind during the second half of the occultation: 15m/s.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Sozzetti_22 Wasp33 HKspec, 200_H+K 5h2m Tell. We had to interrupt integration due to a GCS problem.
Hunt_08 GRB020127 zJspec, 200_H+K 50m We repeat the zJ integration to compensate for slit loss during the bright time observations
La Franca_A12 NGC4686 zJspec, 200_H+K 45m Tell. COMPLETE

11.35 We perform dark and flat calibrations for Sozzetti_22 program
  • ./DoDarks luci.20121207.0001-0005
  • only dark_nor.txt
  • ./Doflat luci.20121207.0006-0015

12.38 The T.O. opens the Dome

01.35 We slew to Sozzetti_22 and point WASP33
  • executeLuciscript.sh spec_acq_science_22NEW.text luci.20121207.0021-0023
  • FILTER entry missing in acq script
  • Images saturated. Flush detector required

01:42 executeLuciscript.sh clean1.txt
  • Aborted after 85 exposures (no images saved)

01:56 executeLuciscript.sh spec_acq_science_22NEW.text
  • Guiding error: could't find star

01.58 loose of collimation. T.O. slews to a nearby bright star to focus

2.00 executeLuciscript.sh spec_acq_science_22NEW.text luci.20121207.0024-0028

02:15 executeLuciscript.sh spec_science_A_22new.text luci.20121207.0029-0073

02:43 Guide problem (gcs)
  • Script interrupted at 2:45

02:49 executeLuciscript.sh spec_acq_science_22NEW.text
  • We are not able to reproduce exactly the same overlap of detector and sky of the previous frames
  • We decide to start a new acquisition

03:18 executeLuciscript.sh spec_acq_science_22NEW.text luci.20121207.0076-0079
  • Comparing the images luci.20121207.0028 and luci.20121207.0079 the shift in position is -2px,+2px respectively on the X and Y axes.

03:21 executeLuciscript.sh spec_science_A_22new.text luci.20121207.0080-0252 *average seeing 0.7"

05.16 Flexure compensation issue due to transit of the target (elevation > 80deg). We stop the spectroscopy, switch off and then on the flexure comp from LUCI software
  • executeLuciscript.sh spec_science_A_22NEW.text luci.20121207.0253-0530
  • average seeing 0.8"

08:24 Stop Wasp33 integration

08:25 We slew to Hunt_08AB
  • executeLuciscript.sh spec_acq GRB020127acquisition_tel.script luci.20121207.0531-0533

08.35 executeLuciscript.sh GRB020127science_zJspec_tel.script luci.20121207.0534-0542

08.49 executeLuciscript.sh GRB020127acquisition.script luci.20121207.0543-0546

09.04 executeLuciscript.sh GRB020127science_zJspec.script luci.20121207.0547-0551
  • average seeing 1.0"

10:09 Slew to La Franca A12
  • executeLuciscript.sh STDA12_1_J.acq luci.20121207.0552-0553
  • Guide lost: Target is too close the moon and guiding star is too faint

10:15 executeLuciscript.sh STDA12_1_200_zJspec.spec luci.20121207.0558-0561

10:26 executeLuciscript.sh OB_LaFranca_A12_acq.txt luci.20121207.0562-0568
  • We decide to modify DIT/NDIT configuration because the telluric is too bright
  • Condensation spots appears on the FOV

10:50 executeLuciscript.sh OB_LaFranca_A12_spe.txt luci.20121207.0569-0576
  • We decide to change offset according with humidity spot on the detector to avoid the dirty area

11:43 We slew to LaFranca_A6
  • executeLuciscript.sh OB_LaFranca_A6_acq.txt luci.20121207.0577
  • Target too close to the Moon. No collimation

12:03 Dome close due to humidity at 93%

-- EleanoraSani - 07 Dec 2012
Topic revision: r7 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback